OpenStudy (anonymous):

what is the 3 therm in the expansion of (3x-2)^5

OpenStudy (anonymous):

(r+1)th term = nCr a^(n-r)b^r for (a+b)^n 3rd term ofd (3x-2)^5 = 5C2 (3x)^(5-2) (-2)^2 = 10 * 27x^3 * 4 = 1080x^3